OpenStudy (antonio_xx2):

A stock gained 2 points on monday lost 5 point on tusday lost 1 point on wednesday gained 4 points on thursday and 6 point on friday a. was the net change for the week positive or negative? b.how much was the gain or loss?

OpenStudy (anonymous):

+2-5-1+4+6 = -4 + 10 = 6

OpenStudy (anonymous):

so a) positive b) 6

OpenStudy (antonio_xx2):

thanks